OpenStudy (anonymous):

Wes has been tracking the population of a town. The population was 3,460 last year and has been increasing by 1.35 times each year. How can Wes develop a function to model this? @Isaiah.Feynman

OpenStudy (isaiah.feynman):

I'm thinking..f(x)=3,450+1.35x. But I'm not very sure.

OpenStudy (anonymous):

f(x) = 3460(1.35)x where 1.35 is the rate of growth f(x) = (3460 • 1.35)x where 1.35 is the rate of growth f(x) = 1.35(3460)v where 3460 is the rate of growth f(x) = 3460(x)1.35 where 3460 is the rate of growth these are the options

OpenStudy (anonymous):

@Isaiah.Feynman

OpenStudy (isaiah.feynman):

First one.

OpenStudy (anonymous):

Thankss <#3

OpenStudy (anonymous):

This wass correct btw :b